# Natural Language Processing RELIES on Linguistics

### v4

@@ 1 Introduction
It is 2025. ChatGPT has been an international sensation for over two years, a focal point in the ongoing public hype, enthusiasm, and concern about “generative AI.” NLP conferences are awash in papers about prompting Large Language Models that have gobbled zillions of words of text and can produce eerily fluent language (in English, anyway). These models have been refined with corrective feedback on conversational outputs, but (so far as we know) no direct inductive biases about grammaticality or compositional meaning. [^1] For example, we did not find any such linguistic biases in the reported Llama 3.1 training procedure, though parts of the training targeted code (Llama Team, 2024). Like BERT and company an NLP research epoch before—or word embeddings an epoch before that—they have elicited both fascination and hand-wringing in scientific circles.
@@
What do we mean by “linguistics”, “NLP”, and “CL”? Before developing our argument, we need to define our terms (while acknowledging that our definitions will not be perfectly crisp, as disciplinary boundaries are inherently fuzzy).
Linguistics is the study of systematicity and variation in human communication, as transmitted via speech, sign, and writing. Linguists may specialize in how knowledge of forms is organized and deployed to symbolize meanings; how linguistic behaviors emerge, drawing on human cognitive and social capacities; and how language varies over time (ontogeny, phylogeny), geography, identity group, etc. Linguistics-adjacent fields of language study and practice include language education, translation, rhetoric, communications, and philosophy of language. Below, the term “linguistics” covers these (non-AI) “language fields”, broadly defined.
Natural Language Processing (NLP) is the field concerned with developing technology for sophisticated computational processing of text, and especially, computational understanding or generation of individual sentences, documents, or conversations (as opposed to drawing inferences about entire collections). [^3] Text mining, information retrieval, and speech processing are studied in separate research communities, with some overlap. These areas, in combination with prototypical NLP applications such as machine translation and summarization, constitute language technologies. The shifting technological landscape may be opening up opportunities for greater unification of these fields (e.g., Chrupała, 2023). Contemporary NLP research prioritizes empirical evaluation of systems on tasks, whether connected with general user applications (like QA, translation, summarization) or more granular and focused on aspects of the language system (like parsing and coreference resolution). In recent years, the dominant concern of NLP has been the design, interpretation, and application of pretrained transformer language models such as BERT and GPT. Pretrained models of considerable scale are often called Large Language Models (LLMs).
@@
  • Low-resource settings: We would struggle more to understand why approaches that work well for English or French might not work well for Swahili or Arapaho. We would not have the knowledge that would allow us to test linguistic inductive biases in neuro-symbolic models for greater accuracy.
@@
In what follows, we give a sampling of the ways linguistics continues to play a role in NLP, organized under the mnemonic “RELIES”: resources (section 3), evaluation (section 4), low-resource settings (section 5), interpretability/explanation (section 6), and study of language (section 7). It is worth emphasizing that these categories and the literature review are meant to be illustrative, not exhaustive; that we are speaking at a macro level; and that we are not claiming that linguistics is the only or necessarily the single most important source of expertise for working with language data and systems. We conclude by discussing pertinent challenges for the NLP community which involve linguistic expertise (section 8).
@@ 2 Background
Today it is not a given that theories and representations from linguistics will form a direct foundation of NLP technologies; it is possible to do research in NLP without traditional training in linguistics. As was the case in the age of the “statistical revolution” of NLP (Johnson, 2009; Uszkoreit, 2009; Church, 2011), deep learning was heralded as a ‘tsunami’ (Manning, 2015) and a way to model language ‘from scratch’, rather than via features reflecting linguistic expertise Collobert et al. (2011). Milestones in the neural era included the introduction of word embeddings Mikolov et al. (2013) and then pretrained language models Howard and Ruder (2018), with the well-known inflection points of ELMo and BERT Peters et al. (2018); Devlin et al. (2019) in 2018/19 and ChatGPT Open AI (2022) in 2022/23. The recent past has seen LLMs with billions of parameters that can be accessed by a wide range of users through a natural language interface (‘prompting’). Do these developments prove the ‘bitter lesson’ Sutton (2019) that language technologies are most effectively developed solely by funneling data on a massive scale into rather general machine learning architectures? What role does linguistics play nowadays in NLP?
We are not the only researchers to reflect on these apparently tectonic shifts. Ignat et al. (2024) collect interesting topics for PhD students, particularly evaluation, and Saphra et al. (2024) hypothesize a cyclical historic model that then would suggest that familiar problems will resurface. We also see that the relationship between Linguistics and NLP (as well as the relationship between Linguistics and AI) has been a topic of discussion for some time (e.g., Lakoff, 1978; Raskin, 1985; Nirenburg, 1986; Baldwin and Kordoni, 2009; Xia, Lewis, and Levin, 2010). In 2011, an entire collection of papers was devoted to the relationship between linguistics and CL (Baldwin and Kordoni, 2011). There it was noted that the field of CL encompasses both scientific and engineering perspectives (Kay, 2011; Johnson, 2011); and that even with machine learning algorithms, long-tail (rare) phenomena celebrated by linguists are likely to challenge systems (Kay, 2011; Levin, 2011; Steedman, 2011), and claims that NLP solutions will work well for any language need to be carefully evaluated in light of linguistic diversity (Bender, 2011) and low-resource conditions in small languages (Bird, 2011).
In what follows, we offer a contemporary take on how linguistics has enduring, practical relevance for NLP research, under the mnemonic “RELIES”.
@@ 3 Resources
The selection and curation of data can be informed by language expertise, even if the data is raw text. We wish for reflection of language diversity, variation among dialects/speakers, genre variation, code-switching, and (if not target of a specific RQ, probably) reduced social biases or problematic content (e.g., hate speech). Sensitivity to linguistic factors is particularly important because of the wide variability of human language and the subjectivity inherent in certain labeling problems, like sentiment and toxic language (Sasidharan Nair, Dinkar, and Abercrombie, 2024; Abercrombie, Hovy, and Prabhakaran, 2023). Specific data selection techniques may involve shallow linguistic heuristics such as frequency counts, or linguistically advanced techniques that increase and validate the diversity and language phenomena coverage of the data Dryer (1989); Rijkhoff et al. (1993). For example, Ravichander, Gardner, and Marasovic (2022) select data with sufficient prepositions and complementizers, and Ponti et al. (2020) inform their data selection with a language typology index Littell et al. (2017). The BabyLM challenge (Warstadt et al., 2023; Hu et al., 2024) selects training and testing data according to cognitive linguistic plausibility, addressing fundamental questions in the learning process of LLMs.
In the annotation process, advanced linguistic knowledge of annotators has been of value, and people skilled at language analysis can ensure meaningful evaluation of NLP systems. For example, Freitag, Grangier, and Caswell (2020) observe that when linguists paraphrase MT references, this increases their quality and diversity; And only professional translators (instead of crowd-workers) lead to correct MT system rankings (Freitag et al., 2021a). For consistent sentiment annotation, Taboada (2016) highlights the importance of understanding linguistic discourse phenomena. Since not all tasks may be suitable for crowdsourcing without extensive experimentation, linguistically trained annotators can help reduce annotation cost (Gillick and Liu, 2010).
Finally, designing and coordinating annotation protocols, especially via crowdsourcing, requires sensitivity to linguistic issues in order to craft guidelines and processes that produce a reasonable level of coherence across the annotations, and to recognize where disagreements can be explained by different perspectives (Ambati, Vogel, and Carbonell, 2012; Plank, 2022; Wein et al., 2023; Prabhakaran et al., 2024).
@@
Two examples of linguistically detailed, resource-oriented frameworks prominent in current NLP research are Universal Dependencies, with morphosyntactic annotations for over 150 languages (UD; Nivre et al., 2016, 2020), and Abstract Meaning Representation, which is a framework for describing semantic graphs that has datasets in several languages (AMR; Banarescu et al., 2013; Wein and Schneider, 2024). UD and AMR are applied linguistic theories (in the sense of Raskin, 1985, p. 275). Research in and with linguistic annotation frameworks continues apace, with numerous recent workshops in the ACL community focused on grammar and semantics (Bonial and Tayyar Madabushi, 2023; Dakota et al., 2023; Bonn and Xue, 2023; Grobol and Tyers, 2023; Rambow and Lareau, 2023). The AMR framework alone is featured in 450 papers, [^4] https://nert-nlp.github.io/AMR-Bibliography/ and numerous applications (Wein and Opitz, 2024; Sadeddine, Opitz, and Suchanek, 2024).
Generally, linguistically-annotated corpora and NLP models contribute to studying questions about language (section 7). However, for some major NLP applications (like MT and QA) in high-resource settings, the future of explicit linguistic representations remains to be seen, with mixed results from neuro-symbolic models (Hamilton et al., 2022; Shwartz, 2023). On the other hand, we consider the utility of such formal representations for evaluating and interpreting models (e.g., Xu et al., 2021; Opitz and Frank, 2022; Fodor, De Deyne, and Suzuki, 2024) to be an exciting path (Evaluation, section 4; Interpretability, section 6).
@@ 4 Evaluation
We illustrate where linguistics is useful to successfully evaluate an NLP system, that is, to describe the degree to which it conforms to our expectations. For tasks that take NL input or generate NL output, it is important to measure correctness and robustness at the language-system level, for which appropriate evaluations (quantitative and qualitative) are informed by linguistic expertise. [^5] A similar sentiment was expressed by Levin (2011) in the statistical NLP era: “language technologists [ought to] understand the object of study, human language…in order to understand where current methods are falling short, we as a field need to understand the data” (p. 19).
@@
Linguistic resources, particularly so-called gold standard annotations, play an important role in the evaluation of NLP technologies. They also contribute to the benchmarking of LLMs and “AI systems” (Tenney et al., 2019a; Pimentel et al., 2020a): Many of the tasks packaged in the Beyond the Imitation Game Benchmark (BIG-bench) (Srivastava et al., 2023) are linguistic tasks, such as bridging resolution and detecting common morphemes. Another kind of benchmark tests linguistic inference from just a few examples in puzzles (Şahin et al., 2020; Chi et al., 2024), a format conducive to low-resource languages without established annotation conventions (Bean et al., 2024).
Linguistic resources are also crucial for system diagnostics (including error analysis), which is a part of evaluation, but does not necessarily have the aim to produce a “global” ranking of AI systems. Instead, diagnostic methods help find more targeted improvement perspectives, with possible relevance even to broader society, since linguistic phenomena are often contingent with issues of societal interest, like gender biases in coreference (Rudinger et al., 2018) or QA (Parrish et al., 2022). [^6] Raji et al. (2021) generally argue for more targeted evaluation methods to track “AI progress.”
Overall, linguistic diagnostics can come in many forms; here we give a small sample. The ‘CheckList’ (Ribeiro et al., 2020) tests NLP systems in various aspects, using linguistic features (negation, part of speech, etc.) to build templates. Song et al. (2022) assess LLMs on Chinese linguistic phenomena, and Parcalabescu et al. (2022) employ linguistic aspects to study (computer) vision and language systems. Automated diagnostics can be achieved, e.g., through targeted measurements on parsed semantic structures (Lo and Wu, 2011; Opitz and Frank, 2021; Fan, Aumiller, and Gertz, 2023, calling for work on the intrinsic NLP task of parsing), or stratifying evaluation data by semantic complexity (Antoine et al., 2024). Then there is a large area of probing systems with linguistic structures Tenney et al. (2019b); Rama, Beinborn, and Eger (2020); Pimentel et al. (2020b); Starace et al. (2023), also tailored at certain branches of linguistics, e.g., psycholinguistics Gauthier et al. (2020); Ettinger (2020). Moore (2009) has examples of challenges that require linguistic knowledge to unpack, such as structural parallelism and translation of “WH” questions.
In sum, linguistics is crucial for diagnostics of NLP technologies. Besides tracking system progress, it fosters interpretability (section 6), and calls for resource creation (section 3).
@@
The combination of varied human evaluation practices with progress towards highly fluent models has led to diminished reliability of human evaluations: Since generated text has become so fluent, it is harder for humans to distinguish their quality. As a result, not all human evaluation studies are useful measures of model output (Clark et al., 2021; Freitag et al., 2021a), necessitating standardization with regard to design and terminology (Howcroft et al., 2020; Belz, Mille, and Howcroft, 2020; van der Lee et al., 2019). Making use of implicit linguistic knowledge held by native speakers (e.g., judging the fluency of system output) or more explicit linguistic training (e.g. assessing the syntactic diversity of system output), can serve a crucial role in reliably assessing the state of the field (Michael et al., 2016; Callison-Burch et al., 2008).
To design approaches for meta-evaluation (assessing and comparing metrics), expertise in linguistics can prove highly useful, since the challenging linguistic phenomena for models and metrics need to be understood in order to be identified. In a WMT metric shared task, Freitag et al. (2021b) find that most metrics struggle to handle complex semantic phenomena and do not appropriately punish the reversal of sentiment or negation; this is established with test sets that assess specified linguistic features (here: sentiment or negation reversal).
@@
Linguistic annotations can help build automatic metrics. Some NLG metrics are built from entailment data [^7] While entailment annotations may come from non-linguists, the notion of textual entailment was inspired by traditions from logic and formal semantics Dagan, Glickman, and Magnini (2006); Bowman et al. (2015). (Chen and Eger, 2023; Xie et al., 2021; Scialom et al., 2021; Steen et al., 2023). Measures for biases, like social biases, have drawn upon the efforts of linguists. For instance, Steen and Markert (2024) propose bias evaluation measures from socio-linguistic lexicons (gender, race) and OntoNotes (Pradhan et al., 2013).

### v5

@@ 1 Introduction
It is 2025. ChatGPT has been an international sensation for over two years, a focal point in the ongoing public hype, enthusiasm, and concern about “generative AI.” NLP conferences are awash in papers about prompting Large Language Models that have gobbled zillions of words of text and can produce eerily fluent language (in English, anyway). These models have been refined with corrective feedback on conversational outputs, but (so far as we know) no direct inductive biases about grammaticality or compositional meaning. [^1] For example, we did not find any such linguistic biases in the reported Llama 3.1 training procedure, though parts of the training targeted code (llama3). Like BERT and company an NLP research epoch before—or word embeddings an epoch before that—Large Language Models have elicited both fascination and hand-wringing in scientific circles.
@@
What do we mean by “linguistics,” “NLP,” and “CL”? Before developing our argument, we need to define our terms (while acknowledging that our definitions will not be perfectly crisp, as disciplinary boundaries are inherently fuzzy).
Linguistics is the study of systematicity and variation in human communication, as transmitted via speech, sign, and writing. Linguists may specialize in how knowledge of forms is organized and deployed to symbolize meanings; how linguistic behaviors emerge, drawing on human cognitive and social capacities; and how language varies over time (ontogeny, phylogeny), geography, identity group, etc. Linguistics-adjacent fields of language study and practice include language education, translation, rhetoric, communications, and philosophy of language. Below, the term “linguistics” covers these (non-AI) “language fields,” broadly defined.
Natural Language Processing (NLP) is the field concerned with developing technology for sophisticated computational processing of text, and especially, computational understanding or generation of individual sentences, documents, or conversations (as opposed to drawing inferences about entire collections). [^3] Text mining, information retrieval, and speech processing are studied in separate research communities, with some overlap. These areas, in combination with prototypical NLP applications such as machine translation and summarization, constitute language technologies. The shifting technological landscape may be opening up opportunities for greater unification of these fields (e.g., chrupala-2023-putting). Contemporary NLP research prioritizes empirical evaluation of systems on tasks, whether connected with general user applications (like QA, translation, summarization) or more granular and focused on aspects of the language system (like parsing and coreference resolution). In recent years, the dominant concern of NLP has been the design, interpretation, and application of pretrained transformer language models such as BERT and GPT. Pretrained models of considerable scale are often called Large Language Models (LLMs).
@@
  • Low-resource settings: We would struggle more to understand why approaches that work well for English or French might not work well for Swahili or Arapaho (due to either the lack of data or the features of the language itself). We would not have the knowledge that would allow us to test linguistic inductive biases in neuro-symbolic models for greater accuracy.
@@
In what follows, we give a sampling of the ways linguistics continues to play a role in NLP, organized under the mnemonic “RELIES:” resources (section˜3), evaluation (section˜4), low-resource settings (section˜5), interpretability/explanation (section˜6), and study of language (section˜7). It is worth emphasizing that these categories and the literature review are meant to be illustrative, not exhaustive; that we are speaking at a macro level; and that we are not claiming that linguistics is the only or necessarily the single most important source of expertise for working with language data and systems. We conclude by discussing pertinent challenges for the NLP community which involve linguistic expertise (section˜8).
@@ 2 Background
Today it is not a given that theories and representations from linguistics will form a direct foundation of NLP technologies; it is possible to do research in NLP without traditional training in linguistics. As was the case in the age of the “statistical revolution” of NLP (johnson-2009-statistical; uszkoreit-2009-linguistics; church2011pendulum), deep learning was heralded as a ‘tsunami’ (manning-15-tsunami) and a way to model language ‘from scratch,’ rather than via features reflecting linguistic expertise collobert-nlp-from-scratch. Milestones in the neural era included the introduction of word embeddings NIPS2013_9aa42b31_word2vec and then pretrained language models howard2018universal, with the well-known inflection points of ELMo and BERT peters2018elmo; devlin-etal-2019-bert in 2018/19 and ChatGPT openai in 2022/23. The recent past has seen LLMs with billions of parameters that can be accessed by a wide range of users through a natural language interface (‘prompting’). Do these developments prove the ‘bitter lesson’ sutton2019bitter that language technologies are most effectively developed solely by funneling data on a massive scale into rather general machine learning architectures? What role does linguistics play nowadays in NLP?
We are not the only researchers to reflect on these apparently tectonic shifts. ignat2023phd collect interesting topics for PhD students, particularly evaluation, and saphra2023first hypothesize a cyclical historic model that then would suggest that familiar problems will resurface. We also see that the relationship between Linguistics and NLP (as well as the relationship between Linguistics and AI) has been a topic of discussion for some time (e.g., lakoff-78; raskin-85; nirenburg-86; ws-eacl-vvv-09; ws-acl-common-ground-10). In 2011, an entire collection of papers was devoted to the relationship between linguistics and CL (baldwin-11). There it was noted that the field of CL encompasses both scientific and engineering perspectives (kay-11; johnson-11); and that even with machine learning algorithms, long-tail (rare) phenomena celebrated by linguists are likely to challenge systems (kay-11; levin-11; steedman-11), and claims that NLP solutions will work well for any language need to be carefully evaluated in light of linguistic diversity (bender-11) and low-resource conditions in small languages (bird-11).
In what follows, we offer a contemporary take on how linguistics has enduring, practical relevance for NLP research, under the mnemonic “RELIES.”
@@ 3 Resources
The selection and curation of data can be informed by language expertise, even if the data is raw text. When developing a dataset, we strive for reflection of language diversity, variation among dialects/speakers, genre variation, code-switching, and (if not target of a specific RQ, probably) reduced social biases or problematic content (e.g., hate speech). Sensitivity to linguistic factors is particularly important because of the wide variability of human language and the subjectivity inherent in certain labeling problems, like sentiment and toxic language (sasidharan-nair-etal-2024-exploring; Abercrombie, Hovy, and Prabhakaran, 2023). Specific data selection techniques may involve shallow linguistic heuristics such as frequency counts, or linguistically advanced techniques that increase and validate the diversity and language phenomena coverage of the data dryer1989large; rijkhoff1993method. For example, ravichander-etal-2022-condaqa select data with sufficient prepositions and complementizers, and ponti-etal-2020-xcopa inform their data selection with a language typology index littell-etal-2017-uriel. The BabyLM challenge (warstadt-etal-2023-findingsbaby; hu2024findingsbaby) selects training and testing data according to cognitive linguistic plausibility, addressing fundamental questions in the learning process of LLMs.
In the annotation process, advanced linguistic knowledge of annotators has been of value, and people skilled at language analysis can ensure meaningful evaluation of NLP systems. For example, freitag-etal-2020-bleu observe that the quality and diversity of MT references improves when they are paraphrased by linguists. Additionally, only professional translators (rather than crowd-workers) lead to correct MT system rankings (freitag_etal_2021_mqm). For consistent sentiment annotation, sentiment highlights the importance of understanding linguistic discourse phenomena. Since not all tasks may be suitable for crowdsourcing without extensive experimentation, linguistically trained annotators can help reduce annotation cost (gillick-liu-2010-non).
Finally, designing and coordinating annotation protocols, especially via crowdsourcing, requires sensitivity to linguistic issues in order to craft guidelines and processes that produce a reasonable level of coherence across the annotations, and to recognize where disagreements can be explained by different perspectives (workflow_crowdsourcing; plank-22; wein-etal-2023-follow; prabhakaran-etal-2024-grasp).
@@
Two examples of linguistically detailed, resource-oriented frameworks prominent in current NLP research are Universal Dependencies, with morphosyntactic annotations for over 150 languages (UD; nivre-etal-2016-universal; nivre-etal-2020-universal), and Abstract Meaning Representation, which is a framework for describing semantic graphs that has datasets in several languages (AMR; banarescu2013abstract; wein_coli). UD and AMR are applied linguistic theories (in the sense of raskin-85, p. 275). Research in and with linguistic annotation frameworks continues apace, with numerous recent workshops in the ACL community focused on grammar and semantics (cxgsnlp-2023-international; tlt-2023-international; dmr-2023-international; udw-2023-universal; depling-2023-international). The AMR framework alone is featured in over 450 papers, [^4] https://nert-nlp.github.io/AMR-Bibliography/ and numerous applications (wein-opitz-2024-survey; sadeddine-etal-2024-survey).
Generally, linguistically-annotated corpora and NLP models contribute to studying questions about language (section˜7). However, for some major NLP applications (like MT and QA) in high-resource settings, the future of explicit linguistic representations remains to be seen, with mixed results from neuro-symbolic models (hamilton2022neuro; shwartz-fruitless-2023). On the other hand, we consider the utility of such formal representations for evaluating and interpreting models (e.g., xu-etal-2021-dynamic; opitz-frank-2022-sbert; fodor2024coling) to be an exciting path (Evaluation, section˜4; Interpretability, section˜6).
@@ 4 Evaluation
We illustrate where linguistics is useful to successfully evaluate an NLP system, that is, to describe the degree to which it conforms to our expectations. For tasks that take language input or generate language output, it is important to measure correctness and robustness at the language-system level, for which appropriate evaluations (quantitative and qualitative) are informed by linguistic expertise. [^5] A similar sentiment was expressed by levin-11 in the statistical NLP era: “language technologists [ought to] understand the object of study, human language…in order to understand where current methods are falling short, we as a field need to understand the data” (p. 19).
@@
Linguistic resources, particularly so-called gold standard annotations, play an important role in the evaluation of NLP technologies. They also contribute to the benchmarking of LLMs and “AI systems” (tenney2018what; pimentel-etal-2020-information): Many of the tasks packaged in the Beyond the Imitation Game Benchmark (BIG-bench) (srivastava2022beyond) are linguistic tasks, such as bridging resolution and detecting common morphemes. Another kind of benchmark tests linguistic inference from just a few examples in puzzles (sahin-etal-2020-puzzling; chi-etal-2024-modeling), a format conducive to low-resource languages without established annotation conventions (bean2024lingoly).
Linguistic resources are also crucial for system diagnostics (including error analysis), which is a part of evaluation, but does not necessarily have the aim to produce a “global” ranking of AI systems. Instead, diagnostic methods help find more targeted improvement perspectives, with possible relevance even to broader society, since linguistic phenomena are often contingent with issues of societal interest, like gender biases in coreference (rudinger-etal-2018-gender) or QA (parrish-etal-2022-bbq).666raji2021ai generally argue for more targeted evaluation methods to track “AI progress.” Overall, linguistic diagnostics can come in many forms; here we give a small sample. The ‘CheckList’ (ribeiro-etal-2020-beyond) tests NLP systems in various aspects, using linguistic features (negation, part of speech, etc.) to build templates. song-etal-2022-sling assess LLMs on Chinese linguistic phenomena, and parcalabescu-etal-2022-valse employ linguistic aspects to study (computer) vision and language systems. moore-2009-computational discusses examples of challenges that require linguistic knowledge to unpack, such as structural parallelism and translation of “WH” questions. Automated diagnostics can be achieved, e.g., through targeted measurements on parsed semantic structures (lo-wu-2011-meant; opitz-frank-2021-towards; fan-etal-2023-evaluating, calling for work on the intrinsic NLP task of parsing), or stratifying evaluation data by semantic complexity (antoine-etal-2024-linguistically). Then there is a large area of probing systems with linguistic structures tenney2018what-probing1; rama-etal-2020-probing; pimentel-etal-2020-information-probing3; starace-etal-2023-probing4, also tailored at certain branches of linguistics, e.g., psycholinguistics gauthier-etal-2020-syntaxgym; ettinger2020bert.
@@
The combination of varied human evaluation practices with progress towards highly fluent models has led to diminished reliability of human evaluations: Since generated text has become so fluent, it is harder for humans to distinguish their quality. As a result, not all human evaluation studies are useful measures of model output (clark-etal-2021-thats; freitag_etal_2021_mqm), necessitating standardization with regard to design and terminology (howcroft-etal-2020-twenty; belz-etal-2020-disentangling; van-der-lee-etal-2019-best). Making use of implicit linguistic knowledge held by native speakers (e.g., judging the fluency of system output) or more explicit linguistic training (e.g. assessing the syntactic diversity of system output), can serve a crucial role in reliably assessing the state of the field (michael-etal-2016-putting; callison-burch-etal-2008-meta).
To design approaches for meta-evaluation (assessing and comparing metrics), expertise in linguistics can prove highly useful, since the challenging linguistic phenomena for models and metrics need to be understood in order to be identified. In a WMT metric shared task, freitag-etal-2021-results find that most metrics struggle to handle complex semantic phenomena and do not appropriately punish the reversal of sentiment or negation; this is established with test sets that assess specified linguistic features (here: sentiment or negation reversal).
@@
Linguistic annotations can help build automatic metrics. Some NLG metrics are built from entailment data [^7] While entailment annotations may come from non-linguists, the notion of textual entailment was inspired by traditions from logic and formal semantics dagan_nli; bowman-etal-2015-large. (10.1162/tacl_a_00576; xie-etal-2021-factual-consistency; scialom-etal-2021-questeval; steen-etal-2023-little). Measures for biases, like social biases, have drawn upon the efforts of linguists. For instance, steen-markert-2024-bias propose bias evaluation measures from socio-linguistic lexicons (gender, race) and OntoNotes (pradhan-etal-2013-towards).